# Near-limit quantum control beyond analytic tractability in many-body spin systems

### v2

@@ Methods
AHT-Derived Performance Indicators
@@
Expanded pulse alphabet and search-space construction
@@
The inclusion of rotations introduces non-Clifford operations, which transform Pauli operators into linear combinations rather than mapping them within the Pauli group. As a result, the compact algebraic rules that underlie conventional AHT-based filtering no longer apply, rendering analytic pruning strategies ineffective (?). This enlarged discrete sequence space represents a practical compromise between expressiveness and computational tractability prior to the availability of a predictive model.
@@
Neural Network-based performance estimate
@@
Data-Driven Search Tree
@@
where is the total node visits, is the visit count for the current node, and is the best score observed. This upper confidence bound strategy prioritizes less-visited nodes while retaining information from previous evaluations through backpropagation. This approach is designed for discrete, combinatorial optimization problems with rugged landscapes and computationally expensive evaluations. The underlying stochastic tree search algorithm builds on a framework previously developed and validated for learning in complex systems under data-scarce conditions (?).
@@
Experimental and Simulation Efficiency

### v3

@@ Appendix C Decoherence Budget and Calibrated Simulator
Appendix D AHT-Derived Performance Indicators
@@
Appendix E Expanded pulse alphabet and search-space construction
@@
The inclusion of rotations introduces non-Clifford operations, which transform Pauli operators into linear combinations rather than mapping them within the Pauli group. As a result, the compact algebraic rules that underlie conventional AHT-based filtering no longer apply, rendering analytic pruning strategies ineffective [13]. This enlarged discrete sequence space represents a practical compromise between expressiveness and computational tractability prior to the availability of a predictive model.
@@ Appendix F Ensemble simulation
Appendix G Neural Network-based performance estimate
@@
Appendix H Data-Driven Search Tree
@@
where is the total node visits, is the visit count for the current node, and is the best score observed. This upper confidence bound strategy prioritizes less-visited nodes while retaining information from previous evaluations through backpropagation. This approach is designed for discrete, combinatorial optimization problems with rugged landscapes and computationally expensive evaluations. The underlying stochastic tree search algorithm builds on a framework previously developed and validated for learning in complex systems under data-scarce conditions [36].
@@
Appendix I Experimental and Simulation Efficiency